OpenStudy (anonymous):

What property is shown by the equation? Column A Column B 1. (34 + 11) + 9 = (11 + 34) + 9 2. 25 + (–15 + 4) = (25 + (–15)) + 4 3. –(–21 + 8) = 21 + (–8) 4. –30 + (–16) = –16 + (–30) A. commutative property B. associative property C. opposite of a sum property

OpenStudy (anonymous):

–30 + (–16) = –16 + (–30)What property is shown by the equation? Column A Column B 1. (34 + 11) + 9 = (11 + 34) + 9 A.commutative property 2. 25 + (–15 + 4) = (25 + (–15)) + 4 B.associative property 3. –(–21 + 8) = 21 + (–8) C.opposite of a sum property 4. –30 + (–16) = –16 + (–30)

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

@nbarrera_t @Drake6 @Deadly_Roses @Azionne @Anna1448 @AaronAndyson

OpenStudy (azionne):

that would be B

OpenStudy (anonymous):

You have to aline them up with the right one

OpenStudy (azionne):

yea i know its B

OpenStudy (anonymous):

1.b?

OpenStudy (texaschic101):

1. associative 2. associative 3. opposite of sum 4. commutative